### v1

# REALM-Bench: A Real-World Planning Benchmark for LLMs and Multi-Agent Systems

Abstract.
This benchmark suite provides a comprehensive evaluation framework for assessing both individual LLMs and multi-agent systems in real-world planning scenarios. The suite encompasses eleven designed problems that progress from basic to highly complex, incorporating key aspects such as multi-agent coordination, inter-agent dependencies, and dynamic environmental disruptions. Each problem can be scaled along three dimensions: the number of parallel planning threads, the complexity of inter-dependencies, and the frequency of unexpected disruptions requiring real-time adaptation. The benchmark includes detailed specifications, evaluation metrics, and baseline implementations using contemporary frameworks like LangGraph, enabling rigorous testing of both single-agent and multi-agent planning capabilities. Through standardized evaluation criteria and scalable complexity, this benchmark aims to drive progress in developing more robust and adaptable AI planning systems for real-world applications.
LLMs, Multi-Agent Systems, Planning Benchmark
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Conf; February; 2025isbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
As large language models (LLMs) continue to advance in reasoning and planning, as demonstrated by OpenAI’s GPT-4o-Task (OpenAI, 2024), DeepSeek’s R1 (DeepSeek-AI et al., 2025), Anthropic’s Claude 3.5 Sonnet (Anthropic, 2024), and Gemini (Team et al., 2023), the research community is increasingly focusing on developing multi-agent systems (MAS) powered by these models. Recent innovations include AutoGen (Wu et al., 2024), CAMEL (Li et al., 2023), CrewAI (Moura, 2024), LangGraph (LangChain AI, 2024), Dspy (Khattab et al., 2023), and XAgent (XAgent Team, 2023), among others. Although individual LLMs demonstrate significant capabilities, their true potential is realized when they collaborate to tackle complex real-world problems (Chang, 2024).
Most AI benchmarks emphasize perception, language understanding, or basic reasoning. However, real-world challenges, such as supply chain management, disaster response, healthcare logistics, and investment strategies, demand coordinated planning and decision-making among specialized agents. There is a pressing need for robust benchmarks that can evaluate the performance of both single-agent systems and MAS in these complex, high-stakes domains.
1.1. The REALM Benchmark Suite
REALM-Bench (Real-world Planning Benchmark for LLMs and Multi-Agent Systems) addresses the need for rigorous evaluation with carefully curated planning challenges. These scenarios are designed to be both tractable, enabling human validation and debugging, and sufficiently complex to push the boundaries of current AI systems. Each challenge requires reasoning and validation over sequential actions parallel processes, resource constraints, and unexpected disruptions (Chang, 2025, 2023b, 2023a; Chen et al., 2024; Huang et al., 2024).
The suite consists of eleven scenarios that progressively increase in complexity across three key dimensions:
  1. 1.
    Parallel Planning Threads: The number of concurrent planning processes that must be coordinated.
  2. 2.
    Inter-Dependencies: The complexity of relationships and constraints between these planning threads.
  3. 3.
    Disruption Frequency and Impact: The rate and severity of unexpected events that require the adaptation of the plan.
Next, we describe how each scenario can be scaled along these three dimensions.
1.2. Benchmark Scalability
While the base versions of each scenario enable detailed analysis and debugging, they can be scaled along the three dimensions defined above: parallel planning threads, inter-dependencies, and disruption frequency and impact.
For example, an urban ride-sharing scenario becomes increasingly complex as the number of vehicles and passengers grows, with interdependent carpooling routes and frequent traffic disruptions necessitating real-time plan adjustments.
This scalability allows AI planning systems to be evaluated under progressively challenging conditions, while still allowing detailed analysis of failure modes in simpler scenarios.
1.3. Availability and Access
The REALM-Bench Suite V1.0 is available on GitHub (Geng and Chang, 2025). In addition, we plan to host competitions and workshops at major AI conferences in 2025 to foster community engagement and further development.
2. Related Benchmark
Planning benchmarks have evolved from testing basic STRIPS-style planning to evaluating increasingly sophisticated planning capabilities. The International Planning Competition (IPC) has been a primary driver of planning benchmarks since 1998, using PDDL to specify domains like BlocksWorld, Logistics, and Rovers (Taitler et al., 2024). While valuable for testing classical planning algorithms, these benchmarks focus on deterministic environments with complete information and lack the dynamic disruptions common in real-world scenarios.
More recent benchmarks, such as the Process Planning Competition (PPC), have shifted toward continuous processes and temporal constraints (Smith and Johnson, 2020). Their manufacturing scenarios include parallel activities and resource dependencies, but the disruptions remain limited to machine breakdowns with known repair distributions. Similarly, the Dynamic Planning Competition introduces environmental changes during plan execution, yet it focuses primarily on path planning and navigation scenarios (Doe and Roe, 2022).
The annual Automated Negotiation Agents Competition (ANAC), established in 2010, has evolved to incorporate planning elements within its supply chain scenarios (Miller and Davis, 2023). However, its scope remains primarily focused on bilateral negotiations rather than comprehensive planning under uncertainty. For example, the 2024–25 competition featured a main challenge titled “Split the Pie,” an artificial yet simplified negotiation scenario where agents divide resources between parties. The supply chain problems in ANAC do not involve contingency planning, resource reallocation, or adaptation to unexpected disruptions.
Specifically for testing LLMs’ planning capabilities, TimeBench (Chu et al., 2023) and TaskBench (Shen et al., 2023) represent two approaches to evaluating AI planning. TimeBench focuses on temporal reasoning by testing systems’ ability to understand time dependencies and scheduling constraints, though it often relies on synthetic scenarios that fail to capture the dynamic nature of real-world temporal relationships, where deadlines shift and durations remain uncertain. TaskBench, on the other hand, evaluates practical task automation and step-by-step planning; it provides valuable insights into an AI system’s ability to decompose complex goals into manageable steps, but its scenarios may oversimplify the challenges of real-world automation, where outcomes are uncertain and processes are deeply interconnected.
This landscape reveals several gaps in existing benchmarks:
  1. 1.
    Limited Disruption Modeling: Most benchmarks treat uncertainties as static probability distributions rather than dynamic, interdependent events that can cascade through systems.
  2. 2.
    Simplified Dependencies: Real-world planning problems involve rich networks of temporal, resource, and causal dependencies that exceed the complexity found in current benchmarks.
  3. 3.
    Restricted Scope: Benchmarks tend to focus on specific subproblems (path planning, task allocation, etc.) rather than end-to-end planning scenarios that combine multiple challenges.
  4. 4.
    Artificial Constraints: Many benchmarks use simplified representations (like PDDL) that cannot capture the nuanced constraints and objectives found in real-world planning problems.
  5. 5.
    Limited Scalability: Few benchmarks allow systematic scaling of complexity along multiple dimensions while maintaining problem tractability for analysis.
  6. 6.
    LLM Specific Challenges: Although LLMs have achieved remarkable successes, the transformer architecture exhibits certain limitations. For example, an attention sink phenomenon can cause certain tokens to be neglected, potentially skewing model predictions (Xiao et al., 2024). Additionally, maximum likelihood training can introduce biases that limit output diversity and quality (Holtzman et al., 2020; Chang, 2023b). Finally, chain-of-thought approaches may suffer from pitfalls such as error propagation and inconsistent reasoning (Brown et al., 2020; Prystawski et al., 2023). A test suite should specifically examine these LLM-related issues.
Our proposed REALM-Bench addresses these limitations by providing scenarios that combine dependencies, dynamic disruptions, and scalability while remaining tractable for systematic evaluation. This allows testing of planning systems under conditions that better reflect the challenges of real-world applications.
3. Benchmark Structure
Problems are categorized into three difficulty levels based on the number of parallel execution threads, the complexity of dependencies, and real-time disruptions.
3.1. Entry Level (1-2 threads)
Problems focusing on basic coordination with limited dependencies:
    Single or dual thread execution
    Basic timing and resource constraints
    Simple disruption scenarios
    Example: Campus tour coordination with one and two groups
3.2. Intermediate (3-4 threads)
Problems requiring significant coordination across multiple execution paths:
    Three to four parallel threads
    Complex timing dependencies
    Resource sharing constraints
    Example: Wedding logistics with multiple vehicles and tasks
3.3. Advanced (5+ threads)
Problems with real-world complexity:
    Five or more parallel threads
    Complex inter-dependencies
    Multiple resource conflicts
    Dynamic disruption scenarios
    Example: Thanksgiving dinner coordination, natural disaster relief, and supply chain management
3.4. Evaluation Metrics
Each problem is evaluated across five key dimensions:
    Planning Quality: Effectiveness of initial plan generation
    Coordination: Management of parallel thread execution
    Adaptation: Response to disruptions and changes
    Resource Management: Resolution of resource conflicts
    Constraint Satisfaction: Maintenance of problem constraints
4. Benchmark Problem Specifications
Table 1. Single Tour Campus Navigation Problem
Metrics: - Total tour time: Minimize while meeting all constraints - Visit coverage: All locations must be visited Locations: Five campus buildings $L=\{S,L,B,A,D\}$ $S$: Student Center - $L$: Library - $B$: Lab Building - $A$: Athletics Center - $D$: Dormitory Travel Times: (minutes) $S\text{-}L:10,S\text{-}B:15,S\text{-}A:20,S\text{-}D:15$ $L\text{-}B:10,L\text{-}A:25,L\text{-}D:20$ $B\text{-}A:15,B\text{-}D:25$ $A\text{-}D:20$ Visit Requirements: - Start time given - Each location requires 30-minute visit - Tour starts/ends at Student Center ($S$) - Group size: 20 people Time Constraints: - Lab Building: Only 9 AM - 4 PM - Library: After 10 AM - Total tour must complete by 5 PM
Table 2. Multi-Group Campus Tour Problem
Groups: - $G_{1}$: 15 people (domestic students) - $G_{2}$: 20 people (international students) Locations: Ten campus buildings $L=\{S,L,B,A,D,C,M,R,H,P\}$ with capacities $cap_{l}$: $cap_{l}=\begin{cases}40&l\in\{S,A\}\\ 30&l\in\{L,D,C\}\\ 25&l\in\{B,M\}\\ 20&l\in\{R,H,P\}\end{cases}$ Constraints: - Tour start time for each group: between 9 AM and 10 AM - Total visitors must not exceed location capacity - Each location requires 30-minute visit - Both tours start/end at Student Center - Complete all tours by 5 PM Additional Requirements: - Lab tours ($B$) only 9 AM - 4 PM - Dining ($C$) must be visited between 11 AM - 2 PM - Library ($L$) after 10 AM
REALM-Bench comprises eleven foundational problem frameworks that systematically evaluate both sequential and reactive planning. Building on the key dimensions introduced earlier (parallel threads, inter-dependencies, and dynamic disruptions), these frameworks progress from straightforward single-thread execution to complex multi-agent scenarios with real-time challenges.
Considerations #1 Problem Complexity:
Each framework can be further scaled to create more challenging variants:
  • @itemi
    Expanding the scale of agents and resources (e.g., from dozens to thousands)
  • @itemi
    Widening geographic distribution (e.g., local to global)
  • @itemi
    Increasing disruption frequency and severity (e.g., isolated events to cascading failures)
  • @itemi
    Introducing uncertainties in execution times and outcomes (e.g., probabilistic durations)
  • @itemi
    Adding hierarchical dependency networks (e.g., sub-networks with internal dependencies)
  • @itemi
    Accounting for agent properties (e.g., atomicity, idempotency)
This extensible design enables researchers to incrementally assess their systems’ capabilities while preserving each scenario’s fundamental planning challenge.
Considerations #2 Implementation:
Three main approaches can be used to tackle these problems:
  • Rudimentary Manual Approach: Hand-craft a directed graph, choose a solver, and input parameter values to obtain a solution. Although direct, this method demands considerable human effort and domain expertise.
  • LLM-Assisted Semi-automation: Use an LLM to suggest algorithms, extract parameters from specifications, and assist with coding. This approach can handle moderate complexity but becomes less feasible at scale.
  • Full Automation: Provide the problem statement to a multi-agent framework (e.g., MACI (Chang, 2025)) that executes end-to-end, including validation and replanning. This approach, combined with human oversight, scales best for complex scenarios with large numbers of nodes and frequent environmental changes.
Table 3. Urban Ride Sharing Problem
Metrics: - On-time performance: No penalty for early arrivals. - Total distance traveled. Locations: Seven locations: $V=\{A,B,C,D,E,F,G\}$, where $G$ is Boston Logan Airport (BOS). Urban locations $A$$F$ are all 10 km of each other, while distances to BOS are 30+ km. $\begin{bmatrix}&A&B&C&D&E&F\\ A-F&10&10&10&10&10&10\\ \rightarrow G&35&33&36&34&32&31\end{bmatrix}$ Travel speed: ($A$$F$) 60 km/h, and ($A$$F\rightarrow G$) 100 km/h. Passengers: Each passenger specifies an arrival time at BOS ($G$). The dispatcher will instruct drivers when to pick up passengers to ensure on-time arrival at BOS. Ride Requests (Desired BOS arrival time given): - $r_{1}$: Pickup at $A$, to $G$ by 08:45     - $r_{2}$: Pickup at $B$, to $G$ by 08:50 - $r_{3}$: Pickup at $C$, to $G$ by 08:55     - $r_{4}$: Pickup at $D$, to $G$ by 09:00 Available Vehicles (Capacity 2 passengers): - $k_{1}$: at $A$, $k_{2}$: at $C$, and $k_{3}$: at $E$ Scheduling Constraints: - The dispatcher determines the pickup times based on a feasible schedule. Pickup times must allow the driver to first reach the passenger location ($A$ $F$) and then drive to $G$ in time.
4.1. P1: Campus Single-Tour Navigation
Problem Statement: A single autonomous agent must navigate a predefined campus environment to complete a sequence of waypoints while minimizing travel time. The scenario assumes a static environment without disruptions.
Problem Specification:
    Environment: A known map with a finite set of locations.
    Goal: Visit all designated waypoints within a given timeframe.
    Constraints: Opening hours of each location, each location at least 30 minutes, and must be completed before 5 PM.
    Optimization Metric: Shortest path (time or distance).
A meta-plan, shown in Table 1, provides the high-level structure and constraints for the problem. This meta-plan serves as input to specialized solvers, such as dynamic programming or Monte Carlo algorithms, which then generate detailed, executable workflows. The process transforms abstract planning requirements into concrete, implementable sequences of actions while respecting all specified constraints and optimization objectives.
4.2. P2: Multi-Group Campus Tour
Problem Statement: Multiple groups of visitors require guided tours in different locations on a university campus, with optimized scheduling of multiple tour guides. This problem shares the same metrics as P1.
Problem Specification:
    Multiple agents (tour guides) must coordinate to serve different groups of visitors.
    Each group has predefined preferences and time constraints.
    Agents must follow non-overlapping paths while minimizing idle time.
    Location visiting hours must be observed.
Details are provided in Table 2.
4.3. P3: Urban Ride-Sharing (URS)
Problem Statement: Optimize real-time ride assignments for multiple vehicles and passengers in an urban environment, balancing efficiency, fuel use, and service quality.
Problem Specification:
    City Map: A graph $G=(V,E)$ where locations $V$ and roads $E$ have distances and travel times.
    Ride Requests: A set of requests $R$, each defined by:
    • Passenger ID, pickup/drop-off locations, time windows.
    Vehicles: A set of available vehicles $K$, each with:
    • Location, battery/fuel level, passenger capacity and speed.
See Table 3 for detailed specifications.
P4: Urban Ride Sharing (URS) with Disruptions
Problem Specification: Details are provided in Table 4, including two disruption scenarios: airport route traffic delay and local road closure. Other possible disruptions could involve cancelation of passengers or late arrivals.
Table 4. Urban Ride-Sharing Reactive Planning Problem
Vehicles: Three vehicles $V=\{v_{1},v_{2},v_{3}\}$ Capacity: 2 passengers each - Initial vehicle location: City center - Operating hours: all day Passengers: Five passengers $P=\{p_{1},p_{2},p_{3},p_{4},p_{5}\}$ $p_{1}$: Airport by 8:30 AM - $p_{2}$: Airport by 9:00 AM - $p_{3}$: Airport by 9:30 AM - $p_{4}$: Airport by 9:45 AM - $p_{5}$: Airport by 9:45 AM Travel Times: - City center to pickup locations: 15-30 minutes - Pickup locations to airport: 45-60 minutes - Between pickup locations: 20-30 minutes Disruptions: - Airport route traffic delay - Certain local road closure Objectives: - Minimize total vehicle travel time - Meet all passenger deadlines
P5: Wedding Reunion
Problem Specification: Table 5 presents a coordinated wedding event travel problem. Several friends arrive at different times and locations before a 3:00 PM wedding photo session. The challenge includes managing two vehicles for airport pick-ups (aimed at those who cannot drive or wish to cut costs) and completing critical errands, such as collecting the wedding gift and retrieving formal attire from the tailor. All activities must be scheduled to ensure that everyone arrives at the wedding venue before the photo time.
This problem introduces more constraints than the URS problems in P3 and P4, and it also lays the groundwork for a more challenging disruption case discussed in P8.
Table 5. Wedding Reunion Logistics Problem
Metrics: - On-time performance: Must arrive at the venue for 3:00 PM photos. Locations: Four locations: $V=\{B,G,T,W\}$, where $B$ is Boston Airport, $G$ is Gift shop, $T$ is Tailor shop, and $W$ is Wedding venue. Travel time: (minutes) ${}B\text{-}G:45,~{}B\text{-}T:30,~{}B\text{-}W:40,~{}G\text{-}T:20,~{}G\text{-}W:25,~{}T\text{-}W:15$. Arrival Times: - Alex: At $B$ at 11:00 AM from Chicago (need a ride) - Jamie: At $B$ at 12:30 PM from Atlanta (need a ride) - Pat: At $W$ at 12:00 PM driving from NYC (has 5-seater car) Required Tasks: - Gift collection from $G$ (after 12:00 PM) - Clothes pickup from $T$ (by 2:00 PM) - Photos at $W$ (3:00 PM sharp) Available Resources: - One car (5-seater) with Pat, available after he is Boston - Local friend Chris (5-seater) available after 1:30 PM at $W$ Scheduling Constraints: - All tasks must complete before 3:00 PM photo time - Gift store opens at 12:00 PM - Tailor closes at 2:00 PM - Two cars must accommodate all transport needs
P6: Thanksgiving Dinner
Consider a Thanksgiving dinner scenario in which a family of five must return home in a Boston suburb for dinner. The problem involves coordinating departure times, managing travel logistics (including possible traffic delays), and ensuring timely arrival. Table 6 formalizes these challenges as a sequential planning problem.
This scenario also lays the groundwork for a more advanced disruption case, which has proven difficult for standalone LLMs, as discussed in P9.
Problem Specification:
Setup:
  • @itemi
    Mom (Sarah) hosts dinner at 6:00 PM in Boston.
  • @itemi
    Family arrivals:
    • Dad (James) from San Francisco, lands at 1:00 PM ET.
    • Sister (Emily) from Chicago lands at 2:30 PM.
    • Brother (Michael) driving from NY arrives at 3:00 PM.
  • @itemi
    Grandma, who is healthy, needs to pick up nearby.
Constraints:
  • @itemi
    James must rent a car post-landing.
  • @itemi
    Emily needs an airport pickup (no alternatives).
  • @itemi
    Turkey requires 4 hours to cook; someone must be home once it’s in the oven for safety.
  • @itemi
    Side dishes need 2 hours of preparation.
  • @itemi
    Travel times:
    • Home to BOS Airport: 1 hour.
    • BOS Airport to Grandma’s: 1 hour.
    • Home to Grandma’s: 30 minutes.
Key Planning Questions:
  1. 1.
    When should cooking start?
  2. 2.
    Who picks up Emily, and when?
  3. 3.
    When and by whom should Grandma be picked up?
Table 6. Thanksgiving Dinner Coordination Problem
Objective: Coordinate family arrivals and dinner preparation for 6:00 PM dinner in Boston Family Members and Arrivals: - Sarah (Mom): Host, at home - James (Dad): Lands at BOS 1:00 PM from SF - Emily (Sister): Lands at BOS 2:30 PM from Chicago - Michael (Brother): Driving, arrives 3:00 PM from NY - Grandma: Needs pickup from suburban Boston Cooking Requirements: - Turkey: 4 hours cooking time - Side dishes: 2 hours preparation - Someone must stay home during cooking Transportation Constraints: - James must rent car after landing - Emily requires airport pickup - Travel times: Home to BOS Airport: 60 min BOS Airport to Grandma’s: 60 min Home to Grandma’s: 30 min Key Requirements: - All family members at home for 6:00 PM dinner - Turkey and sides ready by dinner time - All pickups completed with available drivers - Cooking supervision maintained
P7: Disaster Relief Logistics Problem
Problem Specification: Table 7 summarizes the problem.
Table 7. Disaster Relief Logistics Problem
Network: $G=(V,E)$ where $V$ represents locations and $E$ represents routes Locations ($V$): - Supply nodes: CW (central warehouse), AP (airport, capacity: 5 tons) - Demand nodes: DC1, DC2, DC3 (distribution centers) - Critical nodes: H1, H2 (hospitals), FS (fuel station) Resources: $cap_{truck}$ $=4\text{ tons},\tau_{truck}=120\text{ min}$ $cap_{heli}$ $=1\text{ ton},\tau_{heli}=30\text{ min}$ Demand Requirements: - DCs: food/water delivery by $T_{dc}=20:00$ Hospitals: medicine within $\Delta T_{h}=6$ hours - Fuel Station: refuel by $T_{fs}=12:00$ Critical Deadlines: - Food/Water: All DCs by 8:00 PM - Medicine: Hospitals within 6 hours - Fuel: FS by 12:00 PM - Airport: Clear excess beyond 5 tons immediately Dynamic Disruptions: - Unpredictable donation arrivals - Road blockages requiring rerouting - Emergency hospital demands - Fuel shortage delays Key Planning Requirements: - Resource distribution scheduling - Transportation mode optimization - Delivery prioritization - Airport overflow management - Real-time disruption handling
P8: Wedding Reunion with Disruptions
Problem Extension: This problem extends P5 with road closures and dynamic rerouting.
The disruption scenario becomes more challenging because, when a road closure is announced, the planner must know each vehicle’s current location to determine whether it is affected. Since LLMs are inherently stateless, they cannot keep track of previous scheduling events and thus struggle to adapt the plan in real-time.
P9: Thanksgiving Dinner with Disruptions
This problem extends P6 by introducing flight delays. Specifically, when a flight from SFO to BOS is delayed by $t$ hours, the new arrival time is confirmed at the originally expected arrival time minus the flight’s scheduled duration. Although this early notice provides an opportunity to adjust travel and dinner plans, current LLM-based systems fail to leverage this information in a timely manner, only beginning to react at the original arrival time and missing the window for earlier intervention.
For example, consider James’s flight, which was originally scheduled to arrive in Boston at 1 PM but has been delayed to 4 PM. He learns of this new arrival time at 10 AM EST (the flight’s intended departure from SFO), offering a three-hour lead for adjustments. However, existing LLM-based solutions do not adapt the plan until 1 PM, thus squandering the opportunity to re-optimize the schedule.
P10: Global Supply Chain
Table 8. Summary of the Data Center Construction Problem Statement
Component Details
GPU Procurement & Shipments
Total GPU Target 100,000 units
Vendors NVIDIA: \$15k/unit, 200k units/quarter capacity, 20% maintenance risk (of unit price over 1 year) AMD: \$10k/unit, 150k units/quarter capacity, 50% maintenance risk (of unit price over 1 year)
Order Timing All orders placed on Day 1
Shipment Schedule Quarterly shipments; e.g., a 600,000 unit order results in 3 shipments of 200,000 units at the end of Q1, Q2, and Q3
Payment Terms 50% deposit at order (day 1), 50% at delivery (each quarter)
Natural Disaster Impacts
Risk Probabilities Earthquake: 10% per quarter; Typhoon: 10% per quarter (assessed over entire project)
Delay Impact Cascading 1-month delay on all subsequent shipments if a disaster occurs in a quarter
Cost Impact 30% price increase applied only to the affected quarter’s shipment
Shipment Expedition Impact Expedite one shipment by 1 month and all subsequent shipments are moved up by 1 month (lock stepped)
Cluster Construction & Infrastructure
Cluster Definition Each cluster requires 50,000 GPUs; clusters can be built concurrently once required GPUs are delivered
Networking Infrastructure (per cluster)
Duration 1 months
Cost \$25 million
Dependencies Can begin only after GPUs have arrived, power is ready, and cooling is operational
Power & Cooling (per cluster)
Requirements Power: 150 MW; Cooling: 1M gallons/day
Lead Time Options Regular: 2 months at \$30M per cluster Expedited: 1 months at \$75M per cluster (150% cost premium)
Testing & Certification (per cluster)
Normal Process 2 months, \$15 million
Expedited Process 1 month, \$45 million (150% cost premium)
Expedition Options Expedite power/cooling, networking (50% time reduction, 150% cost premium), and testing independently per cluster
Business Impact of Delays
Delay Costs \$50 million lost revenue + \$10 million additional operating costs per month of delay
Overall Objective
Objective Minimize the end-to-end construction cost (procurement, construction, maintenance, delay costs, expedition premiums) while meeting the 15-month deadline. If the 15-month not met, adding delay costs
Table 8 presents a comprehensive problem in data center GPU deployment that captures the complexity of large-scale infrastructure projects. The objective is to complete a 1 million GPU data center in 15 months while minimizing total costs. The problem encompasses procurement decisions between NVIDIA ($15$k/unit) and AMD ($10$k/unit) GPUs, where each vendor has different maintenance risks (20% vs 50% of unit price over one year) and quarterly shipment capacities.
The construction process is organized around 50,000 GPU clusters, which require coordinated deployment of power, cooling, and networking infrastructure. Each cluster demands significant resources: 150 MW of power capacity and 1 million gallons per day of cooling water. Infrastructure development follows strict dependencies: Power and cooling systems must be operational before networking installation can begin, and each cluster must complete testing before becoming operational.
The problem incorporates real-world complexities such as risks of natural disasters (10% probability per quarter for both earthquakes and typhoons), which can cause cascading delays and cost increases. Management can expedite various components; power / cooling installation can be accelerated from 4 to 2.5 months for a 80% cost premium, while testing can be shortened from 2 to 1 month by doubling the cost.
The financial implications are significant, as each month of delay incurs $60$M in combined revenue loss and additional operating costs. This creates a complex optimization challenge: balancing procurement costs, expedition premiums, and risk mitigation strategies while adhering to physical and temporal constraints in the construction sequence.
P11: Stock Prediction/Forecasting
Consider a stock market prediction scenario where an automated system must forecast future stock prices while integrating multiple data streams and accounting for market dynamics. The problem involves processing real-time data, managing prediction updates, and responding to market events. Table 9 formalizes this as a sequential planning problem, presenting a comprehensive framework for building an adaptive prediction system. The problem statement details the requirements for creating a robust prediction pipeline that can handle real-time market data, maintain high accuracy, and adapt to changing market conditions. A sample workflow demonstrating the system’s architecture and data flow is provided in Table 9.
Market DataNews FeedEconomic DataData CollectionFeature ExtractionModel Training $t_{refresh}=5min$ $t_{process}=5min$ TrainValid?PredictionAccurate?Trading AlertYesNoYesNo
Figure 1. Stock Market Prediction Workflow with Validation Loops. Two decision points validate: 1) model accuracy on testing data, and 2) real-time prediction accuracy for trading notifications.
Problem Specification:
Setup:
  • @itemi
    Objective: Predict a portfolio of stock prices for time $\tau$, enabling actionable insights for risk management, investment decisions, and automated trading.
  • @itemi
    Historical Data:
    • Training period: 2010–2020.
    • Validation period: 2021–2023.
    • Real-time data feed required during operation.
  • @itemi
    Data Sources: Yahoo Finance, SEC EDGAR, news feeds, macroeconomic indicators.
  • @itemi
    Required Features:
    • Technical Indicators: MA, MACD, RSI, Bollinger Bands.
    • Market Context: Sector indices, competitor stock performance, correlation metrics.
    • External Events: Economic reports, policy changes, earnings announcements.
    • Sentiment Analysis: Real-time news, social media, financial disclosures.
Constraints:
  • @itemi
    Processing latency: Under 5 minutes for real-time predictions.
  • @itemi
    Prediction confidence interval: 95% confidence bounds required.
  • @itemi
    Model Dependencies:
    • Data collection before feature extraction.
    • Feature extraction before model training.
    • Model training before prediction generation.
  • @itemi
    Performance Requirements:
    • Mean Absolute Percentage Error (MAPE) $<$ 5% on validation.
    • False alert rate $<$ 1%.
    • System uptime $>$ 99.9% during market hours.
State Space:
  • @itemi
    Processing Nodes ($V$):
    • Data Nodes: DC (data collection), FE (feature extraction).
    • Model Nodes: MT (model training), PG (prediction generation).
    • Integration Nodes: IG (integration), AG (alert generation).
  • @itemi
    Dependencies ($E$):
    • Pipeline: $DC\rightarrow FE\rightarrow MT\rightarrow PG\rightarrow AG$.
    • Feedback: $AG\rightarrow DC$ for adaptive updates.
    • External: Market events affect all nodes.
Market Regime Adaptation:
  • @itemi
    Detect shifts in volatility, liquidity, and trend strength.
  • @itemi
    Adjust prediction models based on regime detection.
  • @itemi
    Implement fallback strategies for extreme market shifts.
Adaptive Refresh Rates:
  • @itemi
    $t_{\text{Drefresh}}$ (data collection) adjusts based on market volatility.
  • @itemi
    $t_{\text{Nrefresh}}$ (news updates) increases during high-impact events.
Decision Module:
  • @itemi
    Trading Strategy: Buy/sell signals, stop losses, risk thresholds.
  • @itemi
    Risk Assessment: Hedging, diversification, leverage constraints.
  • @itemi
    Dynamic Adjustment: Rebalancing frequency, asset weighting, market response.
Backtesting and Iterative Validation:
  • @itemi
    Run models on historical data to evaluate predictive accuracy.
  • @itemi
    Compare performance on different market regimes/conditions.
  • @itemi
    Adjust hyperparameters based on validation results.
Table 9. Stock Market Prediction Problem Framework
Network Structure: $G=(V,E)$ where $V$ represents processing nodes and $E$ represents data flows Processing Nodes ($V$): - Data nodes: DC (data collection), FE (feature extraction) - Model nodes: MT (model training), PG (prediction generation) - Integration nodes: IG (integration), AG (alert generation) Dependencies ($E$): - Technical Analysis: $DC\rightarrow FE\rightarrow MT$ Market Context: External events affect all nodes - Model Chain: $MT\rightarrow PG\rightarrow AG$ Feedback Loop: $AG\rightarrow DC$ for adaptive updates Data Requirements: - Historical: Training (2010–2020), Validation (2021–2023) - Technical Indicators: MA, BB, MACD, RSI, ATR, OBV - Price Data: OHLCV with adjustments - External: Policy, economic indices, news sentiment System Parameters: $t_{refresh}$ $=5\text{ min},\text{conf}_{interval}=95\%$ MAPE $<5\%,\text{uptime}=99.9\%$ Workflow Stages: - Stage 1: Data Collection & Feature Extraction - Stage 2: Model Training & Validation - Stage 3: Prediction Generation - Stage 4: Alert & Decision Support - Stage 5: Iterative Refinement Dynamic Adaptations: - Market regime shifts trigger model updates - News events modify prediction weights - Volatility affects refresh rates - Anomalies initiate retraining cycles Performance Metrics: - Prediction accuracy (RMSE, MAPE) - System responsiveness (latency) - Alert precision (false positive rate) - Model adaptability (regime changes)
4.4. Execution Example: Urban Ride Sharing
For a benchmark suite, we encourage users to devise novel methodologies to solve the problems. However, for this illustrative example of Urban Ride Sharing (URS) (P3), we employ a customized specification language that extends PDDL (Planning Domain Definition Language) and workflow networks to support:
    Dynamic constraints and real-time updates.
    Integration with streaming data sources.
    Explicit representation of uncertainty.
    Temporal and spatial dependencies.
The solver, whether manual or automatic, approaches the planning problem in three steps:
  1. 1.
    Convert the problem statement into a formal specification:
      Key objectives and constraints.
      Required resources and their capabilities.
      Performance metrics and success criteria.
      Temporal and spatial dependencies.
  2. 2.
    Transform the specification into a workflow graph:
      Nodes represent processing stages, decision points, or actions.
      Edges capture dependencies, data flow, and execution sequence.
      Agents are assigned to both nodes and edges.
      Concrete parameters and thresholds are specified.
  3. 3.
    Select and apply solving algorithms:
      Test multiple solution approaches (e.g., dynamic programming, Monte Carlo).
      Evaluate solutions against specified metrics.
      Select and validate the best-performing solution.
      Present results with performance analyses.
We walk through these three steps to solve the URS problem.
4.4.1. URS Problem Specification
Already presented in Table 3.
4.4.2. URS Workflow
Figure 2 shows the workflow representation of the URS problem.
ABCDEFG (BOS)$r_{1}$ (8:45)$r_{2}$ (8:50)$r_{3}$ (8:55)$r_{4}$ (9:00)$k_{1}$ [$k_{2}$ [2]$k_{3}$ [2]
Figure 2. Consider a network $G=(V,E)$ with urban travel times $\tau_{ij}=10$ minutes and airport routes $\tau_{iG}=\{19,\ldots,22\}$ minutes. Schedule three vehicles $k_{i}$ (capacity $c_{k}=2$) to deliver four passengers $r_{i}$ to airport during $[8:45,9:00]$.
[…]
URS
4.4.3. URS Results
Figure 3 presents an optimal solution with a total travel distance of 87 km, outperforming both GPT-4o-Task and DeepSeek R1 (Figure 4), which require 123 km. This represents a significant improvement of 41.37%.
Extended Materials
We also used LangGraph to implement both P3 and P4, as presented in Appendix A. Furthermore, we provide a solution to P11 for the prediction of the stock value. These sample implementations validate the completeness of the problem statements in this benchmark.
[…]
Figure 3. This solution routes showing optimal vehicle assignments. Vehicle k1 (blue) starts at A and collects passengers r1 and r2, while k2 (orange) starts at C and serves r3 and r4. Both vehicles arrive at BOS at 8:30, meeting all passenger arrival deadlines. Total distance traveled = 87 km (Optimal).
[…]
This figure illustrates the optimal vehicle assignments for the ALAS solution routes, showing vehicles k1 and k2 serving passengers and meeting deadlines.
[…]
Figure 4. GPT-4o-Task and DeepSeek R1 both schedule three routes. Vehicle k1 picks up r1 from A at 7:40, then r2 from B at 7:50. Vehicle k2 picks up r3 from C at 7:55. Vehicle k3 must first drive from E to D to pick up r4 at 8:00. All meet deadlines. Total travel distance = 123 km.
[…]
Infeasible Schedule
5. Conclusion
REAL-Bench represents a significant step toward systematically evaluating AI systems’ capabilities in real-world planning scenarios. By providing 11 carefully designed problems that progress in complexity, the benchmark enables researchers to:
    Assess planning capabilities in multiple dimensions of difficulty.
    Test system performance on real-world planning challenges.
    Evaluate handling of unexpected interruptions and adaptations.
    Compare different approaches using standardized metrics.
The benchmarks are designed to be both tractable for systematic evaluation and challenging for current systems. Each problem can be scaled along multiple dimensions, including the number of parallel threads, complexity of dependencies, and frequency of disruptions, allowing researchers to progressively stress-test their systems. Inclusion of validation metrics and baseline implementations facilitates meaningful comparisons between different approaches.
Looking ahead, we envision this benchmark suite evolving with community contributions and feedback. Future extensions might include more complex scenarios, additional evaluation metrics, and expanded validation tools. For instance, in many workflows, transaction properties must be preserved:
    Atomicity: An operation either completes entirely or not at all, with no partial execution state (e.g., a ride-sharing trip must either complete fully or be canceled entirely).
    Idempotency: Multiple identical requests produce the same outcome as a single request, preventing duplicate actions (e.g., multiple identical order submissions should not result in multiple orders).
Most importantly, by providing a common framework for evaluating planning capabilities of both individual LLMs and multi-agent systems, we hope to accelerate progress toward more robust and capable AI planning systems that can handle real-world complexity and uncertainty. The REALM benchmark suite, along with detailed documentation and baseline implementations, will be available as an open source resource after the peer review process.
Appendices
In these Appendices, we provide sample implementations of selected problems, illustrating that each problem specification is implementable and can produce feasible solutions. This serves as a verification of completeness for the problem definitions. We encourage readers to review the code and use it as a reference for designing improved solutions to these challenges.
The three problems selected are P3, P4, and P11, representing sequential planning, reactive planning, and the most complex planning scenario, respectively.
Appendix A P3 & P4: Urban Ride Sharing Sample Implementation
This appendix presents an implementation of P3 and P4: Urban Ride Sharing without and with interrupts, using LangGraph (LangChain AI, 2024).
A.1. Agentic Workflow Formulation
In this first stage, we define agents to manage the nodes of the workflow, including data collection, route planning, vehicle dispatch, traffic adjustment, monitoring and alert, and logging agents. At the end, we use the >> syntax to specify dependencies among agents. The transition from problem specifications to workflow formulation is handled automatically by MACI (Chang, 2025) in LangGraph.
Listing 1: Collaborative Agents and Prompts
# ---- Data Collection Agent ---- #
DC_Agent = Agent(
name=‘‘Data Collection Agent’’,
backstory=‘‘You collect basic traffic data, road closure updates, and estimated travel times between locations in Bay Area suburb.’’,
task_description=‘‘Retrieve traffic conditions, road closures, and estimated travel durations for all routes involved in passenger transport.’’,
task_expected_output=‘‘Structured travel time data, including:
1) CityMap: A graph G = (V, E), where the locations V and roads E have distances and travel times.
2) Ride Requests: A set of requests R, each defined by: PassengerID, pickup/drop-off locations, and time windows.
3) Vehicles: A set of available vehicles K, each with location, battery/fuel level, passenger capacity, and speed.’’)
# ---- Route Planning Agent ---- #
RP_Agent = Agent(
name=‘‘Route Planning Agent’’,
backstory=‘‘You determine optimal routes for vehicles to minimize total travel time while ensuring all passengers arrive on time.’’,
task_description=‘‘Use traffic data and constraints to compute the best routes for each vehicle, ensuring on-time airport arrivals.’’,
task_expected_output=‘‘Optimized vehicle assignments and travel routes.’’)
# ---- Vehicle Dispatch Agent ---- #
VD_Agent = Agent(
name=‘‘Vehicle Dispatch Agent’’,
backstory=‘‘You assign passengers to vehicles and ensure each vehicle follows the optimal planned route,’’,
task_description=‘‘Assign passengers to vehicles based on capacity constraints and route efficiency,’’
task_expected_output=‘‘Vehicle assignment list and dispatch schedule.’’)
# ---- Traffic Adjustment Agent (e.g. disruptions, special cases)---- #
TA_Agent = Agent(
name=‘‘Traffic Adjustment Agent’’,
backstory=‘‘You monitor live traffic updates and adjust vehicle routes dynamically in case of delays in Bay Area suburb.’’,
task_description=‘‘Recompute routes in real time when disruptions occur (traffic, road closures), ensuring minimal delays.’’,
task_expected_output=‘‘Updated travel plans for affected vehicles.’’)
# ---- Monitoring & Alert Agent ---- #
MA_Agent = Agent(
name=‘‘Monitoring & Alert Agent’’,
backstory=‘‘You track vehicle movements and notify if there are risks of missing passenger deadlines in Bay Area suburb.’’,
task_description=‘‘Send alerts for potential delays and recommend contingency plans.’’,
task_expected_output=‘‘Timely notifications for alternative route adjustments or emergency responses.’’)
Writer_agent = Agent(
name=‘‘Writer Agent’’,
backstory=‘‘You are a language model specialized in writing text into .json files’’,
task_description=‘‘Write the json response into ./p3_output.json’’,
task_expected_output=‘‘A .json file containing the given string,
tools=write_str_to_txt’’)
# ---- Define Dependencies (With Disruption) ---- #
DC_Agent >> RP_Agent >> VD_Agent >> TA_Agent >> MA_Agent >> Writer_agent
# ---- Define Dependencies (Without Disruption) ---- #
DC_Agent >> RP_Agent >> VD_Agent >> MA_Agent >> Writer_agent
A.2. P3: Execute Meta Plan without Disruption
Now that the meta-plan workflow has been constructed, the second step involves providing real data for workflow execution. The following code snippet illustrates how vehicle and passenger locations are specified, followed by the corresponding agent executions.
Listing 2: Sample JSON Output from Agentic Workflow
RUNNING AGENT: Data Collection Agent
[…]
- V1: Downtown
- V2: Golden Gate Park
- V3: Fisherman Wharf
- V4: Union Square
- V5: San Francisco Airport (SFO)
[…]
- E1: Downtown to Golden Gate Park, Distance: 5 miles, Travel Time: 20 minutes
- E2: Golden Gate Park to Fisherman Wharf, Distance: 6 miles, Travel Time: 25 minutes
- E3: Fisherman Wharf to Union Square, Distance: 2 miles, Travel Time: 10 minutes
- E4: Union Square to SFO, Distance: 15 miles, Travel Time: 30 minutes
- E5: Downtown to SFO, Distance: 13 miles, Travel Time: 35 minutes
[…]
- Request 1:
- PassengerID: P001
- Pickup: Downtown
- Drop-off: Union Square
- Time Window: 9:00 AM - 9:30 AM
- Request 2:
- PassengerID: P002
- Pickup: Golden Gate Park
- Drop-off: SFO
- Time Window: 10:00 AM - 11:00 AM
- Request 3:
- PassengerID: P003
- Pickup: Fisherman Wharf
- Drop-off: Downtown
- Time Window: 1:00 PM - 1:30 PM
[…]
- Vehicle 1:
- Location: Downtown
- Battery/Fuel Level: 80%
- Passenger Capacity: 4
- Speed: 30 mph
- Vehicle 2:
- Location: Union Square
- Battery/Fuel Level: 60%
- Passenger Capacity: 6
- Speed: 25 mph
- Vehicle 3:
- Location: SFO
- Battery/Fuel Level: 90%
- Passenger Capacity: 3
- Speed: 35 mph
This structured data provides an overview of the current traffic conditions, road closures, and estimated travel times for passenger transport in SF.
RUNNING AGENT: Route Planning Agent
Optimized vehicle assignments and travel routes:
[…]
- Vehicle 1:
- Assigned to Request 1
- Route: Downtown (V1) -> Union Square (V4)
- Estimated Travel Time: 20 minutes (E1 + E3)
- Departure Time: 9:00 AM
- Arrival Time: 9:20 AM
- Vehicle 2:
- Assigned to Request 2
- Route: Union Square (V4) -> Golden Gate Park (V2) -> SFO (V5)
- Estimated Travel Time: 55 minutes (E3 + E2 + E4)
- Departure Time: 10:00 AM
- Arrival Time: 10:55 AM
- Vehicle 3:
- Assigned to Request 3
- Route: Fisherman Wharf (V3) -> Downtown (V1)
- Estimated Travel Time: 10 minutes (E3)
- Departure Time: 1:00 PM
- Arrival Time: 1:10 PM
[…]
- Request 1:
- Vehicle 1 will pick up Passenger P001 at Downtown and drop off at Union Square.
- Route: V1 -> V4
- Request 2:
- Vehicle 2 will pick up Passenger P002 at Golden Gate Park and drop off at SFO.
- Route: V4 -> V2 -> V5
- Request 3:
- Vehicle 3 will pick up Passenger P003 at Fisherman Wharf and drop off at Downtown.
RUNNING AGENT: Vehicle Dispatch Agent
Vehicle Assignment List and Dispatch Schedule:
[…]
- Vehicle 1:
- Assigned to Request 1
- Passengers: P001
- Route: Downtown (V1) -> Union Square (V4)
- Estimated Travel Time: 20 minutes
- Departure Time: ‘‘9:00 AM’’
- Arrival Time: ‘‘9:20 AM’’
- Vehicle 2:
- Assigned to Request 2
- Passengers: P002
- Route: Union Square (V4) -> Golden Gate Park (V2) -> SFO (V5)
- Estimated Travel Time: 55 minutes
- Departure Time: ‘‘10:00 AM’’
- Arrival Time: ‘‘10:55 AM’’
- Vehicle 3:
- Assigned to Request 3
- Passengers: P003
- Route: Fisherman Wharf (V3) -> Downtown (V1)
- Estimated Travel Time: 10 minutes
- Departure Time: ‘‘1:00 PM’’
- Arrival Time: ‘‘1:10 PM’’
[…]
- 9:00 AM: Vehicle 1 departs from Downtown with Passenger P001.
- 9:20 AM: Vehicle 1 arrives at Union Square.
- 10:00 AM: Vehicle 2 departs from Union Square with Passenger P002.
- 10:55 AM: Vehicle 2 arrives at SFO.
- 1:00 PM: Vehicle 3 departs from Fisherman Wharf with Passenger P003.
- 1:10 PM: Vehicle 3 arrives at Downtown.
These assignments and schedules ensure efficient use of vehicles and timely arrival of all passengers.
RUNNING AGENT: Monitoring & Alert Agent
[…]
1) Vehicle 1:
- Current Status: On schedule.
- Recommendation: No action needed as the vehicle is expected to arrive on time.
2) Vehicle 2:
- Current Status: Potential delay risk due to high traffic expected around Golden Gate Park.
- Recommendation: Consider adjusting the route to avoid traffic congestion. Possible alternative: Take the route via Sunset Blvd to bypass heavy traffic areas. Notify the passenger of the potential delay and the alternative route.
3) Vehicle 3:
- Current Status: On schedule.
- Recommendation: No action needed as the vehicle is expected to arrive on time.
[…]
Emergency Response: If Vehicle 2 faces unexpected delays despite the alternative route, prepare a standby vehicle for immediate dispatch from a nearby location to ensure Passenger P002 reaches SFO on time.
RUNNING AGENT: Writer Agent
Thought: I need to write the given context string into a JSON file named ‘‘./p3_output.json’’.
Using Tool: write_str_to_txt
Tool call dict: {omitted}
Data successfully written to ‘‘./p3_output.json’’
Tool result: None
Observations: {0: None}
The context string has been successfully written into ‘‘./p3_output.json’’ as a JSON file.
A.3. P4: Execute Meta Plan with Disruption
This section depicts the execution of the meta-plan (workflow) under traffic disruption. In this scenario, the system dynamically identifies traffic delays and integrates real-time updates into the planning process. Each agent plays a crucial role: the data collection agent continuously monitors traffic conditions; the route planning agent adjusts travel routes based on congestion data; and the dispatch agent ensures that vehicle assignments are optimized despite delays. The workflow demonstrates how coordinated agent interventions, including dynamic rerouting and schedule adjustments, effectively mitigate the impact of heavy traffic, maintaining timely transportation, and enhancing overall operational efficiency.
Listing 3: Sample JSON Output from Agent Pipeline
RUNNING AGENT: Traffic Adjustment Agent
[…]
- Start at Union Square
- Travel to Downtown SF (2 miles, 15 minutes)
- Detour via I-280 South instead of the usual route to SF Airport
- Estimated travel time via detour: 30 minutes
[…]
- Start at Fisherman Wharf
- Detour via Bay Street to Van Ness Avenue, then proceed to Golden Gate Bridge
- Estimated travel time via detour: 15 minutes
[…]
Appendix B P11: Stock Prediction Sample Implementation on LangGraph
This appendix presents an implementation of the P11 Stock Prediction problem using LangGraph (LangChain AI, 2024). The implementation demonstrates how AI-driven planning can be applied to financial forecasting by leveraging historical data, technical indicators, and real-time market updates.
B.1. Agentic Workflow Formulation
In this first stage, we define agents to manage the nodes of the workflow, including data collection, feature extraction, model training, integration, and alert. In the end, we use the >> syntax to specify dependencies among agents.
Listing 4: Prompt and Collaborative Agent Pipeline
from src.multi_agent.crew import Crew
with Crew() as crew:
# ---- Data Collection Agent ---- #
DC_Agent = Agent(
name=‘‘Data Collection Agent’’
backstory=‘‘You collect 5 stocks of historical data OHLCV from S&P 500 Kaggle Dataset at https://www.kaggle.com/datasets/andrewmvd/sp-500-stocks, and real-time stock data including external economic indicators.’’,
task_description=‘‘Retrieve stock data from 2023-10-01 to 2023-12-31.’’,
task_expected_output=‘‘Formatted OHLCV data and external indicators, of 5 stocks.’’,
Writer_agent = Agent(
name=‘‘Writer Agent’’,
backstory=‘‘You are a language model specialised in writing text into .json files’’,
task_description=‘‘Write the json response into ./tool_agent_example.json’’,
task_expected_output=‘‘A .json file containing the given string’’,
tools=write_str_to_txt,
# ---- Feature Extraction Agent ---- #
FE_Agent = Agent(
name=‘‘Feature Extraction Agent’’,
backstory=‘‘You compute technical indicators from the collected 5 stock data.’’,
task_description=‘‘Calculate MA, MACD, RSI, Bollinger Bands, and other technical indicators of 5 stocks.’’,
task_expected_output=‘‘Feature matrix of OHLCV, external indicators, MA, MACD, RSI, Bollinger Bands, of 5 stocks.’’
# ---- Model Training Agent ---- #
MT_Agent = Agent(
name=‘‘Model Training Agent’’,
backstory=‘‘You train stock prediction models using historical data (from 2023-10-01 to 2023-11-30) and validate with (from 2023-12-1 to 2024-12-31).’’,
task_description=‘‘use financial math and machine learning models and optimize hyperparameters.’’,
task_expected_output=‘‘Prepare the input Feature matrix, with 3 financial math models and 3 machine learning models ready for inference, of 5 stocks.’’
# ---- Prediction Generation Agent ---- #
PG_Agent = Agent(
name=‘‘Prediction Generation Agent’’,
backstory=‘‘You generate stock price predictions using the trained model.’’,
task_description=‘‘Predict next-day and intra-day stock prices.’’,
task_expected_output=‘‘Based on input and the 6 models, predict Stock price from 2024-1-1 to 2024-1-7, with confidence intervals, of 5 stocks.’’
# ---- Integration Agent ---- #
IG_Agent = Agent(
name=‘‘Integration Agent’’,
backstory=‘‘You combine model predictions with external news sentiment and macroeconomic indicators.’’,
task_description=‘‘Adjust predictions based on news impact and market conditions on year of 2023.’’,
task_expected_output=‘‘Refined Stock price from 2024-1-1 to 2024-1-7, with adjusted confidence intervals, of 5 stocks.’’
# ---- Alert Generation Agent ---- #
AG_Agent = Agent(
name=‘‘Alert Generation Agent’’,
backstory=‘‘You issue buy/sell/hold signals based on predictions and risk assessment of the model’’, task_description=‘‘Generate trading alerts and confidence scores’’,
task_expected_output=‘‘Based on the 6 models, provide actionable trade signals from 2024-1-1 to 2024-1-7, of 5 stocks.’’
# ---- Define Dependencies ---- #
DC_Agent >> Writer_agent >> FE_Agent >> MT_Agent # Technical Analysis Flow
MT_Agent >> PG_Agent >> AG_Agent # Model Chain
#AG_Agent >> DC_Agent # Feedback Loop (adaptive updates)
crew.run()
B.2. Workflow Execution With Historical Data for Prediction
Following the dependency graph presented above (also see Figure 1), the agents are executed in sequence: starting with data collection, then feature extraction, followed by model training, and finally prediction. For brevity, while data for five stocks is provided, we illustrate the execution flow using only APPL.
Listing 5: Sample JSON Output from Data Collection
RUNNING AGENT: Data Collection Agent
json
‘‘stocks’’: [
‘‘symbol’’: ‘‘AAPL’’,
‘‘data’’: [
date: 2023-10-01,
open: 150.00,
high: 152.00,
low: 149.50,
close: 151.00,
volume: 75000000
// Additional daily data up to 2023-12-31
‘‘external_indicators’’: {
interest_rate: 5.25,
inflation_rate: 3.0,
gdp_growth: 2.5
‘‘symbol’’: ‘‘MSFT’’,
‘‘data’’: [
date: 2023-10-01,
open: 280.00,
high: 285.00,
low: 279.00,
close: 282.00,
volume: 45000000
// Additional daily data up to 2023-12-31
‘‘external_indicators’’: {
interest_rate: 5.25,
inflation_rate: 3.0,
gdp_growth: 2.5
‘‘symbol’’: ‘‘GOOGL’’,
‘‘data’’: [
date: 2023-10-01,
open: 135.00,
high: 137.00,
low: 134.50,
close: 136.00,
volume: 32000000
// Additional daily data up to 2023-12-31
‘‘external_indicators’’: {
interest_rate: 5.25,
inflation_rate: 3.0,
gdp_growth: 2.5
‘‘symbol’’: ‘‘AMZN’’,
‘‘data’’: [
date: 2023-10-01,
open: 120.00,
high: 122.50,
low: 119.00,
close: 121.50,
volume: 60000000
// Additional daily data up to 2023-12-31
‘‘external_indicators’’: {
interest_rate: 5.25,
inflation_rate: 3.0,
gdp_growth: 2.5
‘‘symbol’’: ‘‘TSLA’’,
‘‘data’’: [
date: 2023-10-01,
open: 250.00,
high: 255.00,
low: 248.00,
close: 252.00,
volume: 50000000
// Additional daily data up to 2023-12-31
‘‘external_indicators’’: {
interest_rate: 5.25,
inflation_rate: 3.0,
gdp_growth: 2.5
RUNNING AGENT: Feature Extraction Agent
To complete the task of calculating technical indicators for five stocks, perform the following computations for each stock:
Moving Average (MA): Compute the simple moving average over a specified period (e.g., 20 days).
Moving Average Convergence Divergence (MACD): Compute the MACD line, signal line, and histogram.
Relative Strength Index (RSI):} Compute the RSI over a specified period (e.g., 14 days).
Bollinger Bands: Compute the upper band, lower band, and the middle band (the moving average).
Other Technical Indicators: Optionally, compute additional indicators such as Average True Range (ATR), Stochastic Oscillator, etc., based on the available data.
RUNNING AGENT: Model Training Agent
To prepare the input feature matrix with financial math models and machine learning models ready for inference, we will follow these steps:
1. Feature Matrix Preparation:
- Collect historical stock data for the 5 selected stocks, including the following columns: Date, Open, High, Low, Close, and Volume.
- Calculate the necessary financial math models (technical indicators) for each stock:
- Moving Average (MA): Compute the simple moving average over a 20-day period.
- Moving Average Convergence Divergence (MACD): Calculate the MACD line, signal line, and histogram.
- Relative Strength Index (RSI): Calculate the RSI over a 14-day period.
- Bollinger Bands: Calculate the upper band, middle band (moving average), and lower band.
2. Machine Learning Models:
- Train three machine learning models using the historical data and calculated features. The models could be:
- Linear Regression: For predicting future stock prices based on the features.
- Random Forest Regressor: To capture non-linear relationships and interactions between features.
- Support Vector Machine (SVM): For robust predictions with a focus on margin maximization.
- Optimize hyperparameters for each model using techniques such as Grid Search or Random Search to enhance model performance.
3. Inference Preparation:
- Prepare the models for inference by saving the trained models and their hyperparameters.
- Ensure the feature matrix is structured and aligned by date, ready for input into the models for prediction.
4. Output Format:
- Compile the results into a structured data format such as a DataFrame or JSON, including all calculated features and ready-to-use models for each of the 5 stocks.
By following these steps, we ensure that the feature matrix is comprehensive and that the machine learning models are optimized and ready for inference on the selected stocks.
RUNNING AGENT: Prediction Generation Agent
To predict the stock prices from January 1, 2024, to January 7, 2024, for the five selected stocks, we utilize the prepared feature matrix and the trained machine learning models as described in the context. Here is the structured prediction output of one of the five stocks, including confidence intervals:
json
‘‘predictions’’: [
‘‘stock’’: ‘‘APPL’’,
‘‘predictions’’: [
{date: 2024-01-01, predicted_price: 150.25, confidence_interval: [148.00, 152.50], confidence: 0.80},
{date: 2024-01-02, predicted_price: 151.00, confidence_interval: [148.75, 153.25], confidence: 0.80},
{date: 2024-01-03, predicted_price: 152.50, confidence_interval: [150.00, 155.00], confidence: 0.88},
{date: 2024-01-04, predicted_price: 153.75, confidence_interval: [151.50, 156.00], confidence: 0.82},
{date: 2024-01-05, predicted_price: 151.50, confidence_interval: [150.25, 153.75], confidence: 0.90},
{date: 2024-01-06, predicted_price: 155.00, confidence_interval: [152.75, 157.25], confidence: 0.83},
{date: 2024-01-07, predicted_price: 156.25, confidence_interval: [154.00, 158.50], confidence: 0.87}
This output provides the predicted stock prices for each day from January 1 to January 7, 2024, for stock APPL, along with their respective confidence intervals. These predictions are based on the trained models and the prepared feature matrix.
B.3. Output: Alert Generation
Based on the problem specification in Section 4, when the prediction confidence is $\geq 85\%$, a recommended trading action is sent to the subscriber as an alert.
Based on the above prediction, on 1/3, 1/5, and 1/7 the prediction confidence exceeds the threshold and the stock’s moving direction is up, down, and up, respectively. Therefore, the alerts issued before the opening of the market on those three days are: BUY, SELL, and STRONG BUY, respectively.